# The Writing is on the Wall:

### v2

@@ Abstract
Although rollups have attracted significant attention, there is limited empirical research on their performance under high load. To address this, we present a data-driven analysis of the transaction surge in late 2023 and early 2024, attributed to inscriptions—a method for recording data on the blockchain. Initially introduced on Bitcoin, inscriptions enable the representation of NFTs or ERC-20-like tokens without smart contracts, and have since expanded to other blockchains. This paper examines inscription-related transactions on Ethereum and major EVM-compatible rollups, assessing their impact on scalability during transaction surges. Our results show that, on certain days, inscriptions accounted nearly 90% of transactions on Arbitrum and ZKsync Era, while 53% on Ethereum, with 99% of these inscriptions involving meme coin minting. Furthermore, we show that ZKsync and Arbitrum saw lower median gas fees during these surges. ZKsync Era, a ZK-rollup, showed a greater fee reduction than the optimistic rollups studied—Arbitrum, Base, and Optimism.
@@ 1 Introduction
Since the introduction of Bitcoin [25] and later Ethereum [42], blockchains have seen increased adoption [5, 29]. This growth has led to the development of Decentralized Applicationss (DApps), including Decentralized Exchanges (DEXs) [3, 13, 19], lending protocols [28, 31], Non-Fungible Tokens (NFTs)for art [14], and other applications such as supply chain management [30, 32], and decentralized governance [3, 19, 22, 23].
In distributed systems, including blockchains, a fundamental trade-off has been observed; a blockchain can only optimize two out of the following three essential factors: security, decentralization, and scalability [41]; this is the “blockchain trilemma”, a term attributed to V. Buterin [41]. Traditionally, underlying blockchain solutions known as Layer-1 (L1)such as Bitcoin and Ethereum have prioritized decentralization and security at the expense of scalability. To address this challenge, Layer-2 (L2)solutions, built atop an L1, have been proposed. L2 solutions aim to improve scalability and efficiency while relying on the L1blockchain to heritage security. Consequently, a variety of Ethereum Virtual Machine (EVM)-compatible chains have emerged, focusing on scalability without compromising decentralization and security. This is facilitated by executing transactions off-chain and subsequently storing results or state transitions on an L1blockchain.
However, many of these L2scaling solutions were not yet fully stress-tested until late 2023, when a spike in transactions occurred: most L2 EVM-compatible chains experienced their all-year heights of transaction volumes in November and December 2023. This sudden transaction spike was caused by a boom in what is known as inscriptions [1]. Inscriptions originated from a new idea that surged in the Bitcoin blockchain — ordinals [40]. Ordinals were first introduced in March 2023 by an anonymous developer known as Domo [6]. In short, this new system allows users to assign unique identifiers to individual satoshis, the smallest unit of Bitcoin, enabling each satoshi to be tracked and traded separately, based on their ordinal number. This concept adds a new dimension to Bitcoin transactions by introducing the ability to treat each satoshi as a NFT, enabling new applications such as collectible assets, verifiable digital scarcity, and more complex smart contract interactions on the Bitcoin network. This led to a notable surge in the number of transactions in mid-2023 due to a Fear-Of-Missing-Out (FOMO)effect, which subsequently caused an increase in transaction fees for all users [39, 40].
Subsequently, the concept of ordinals was extended to other blockchains, such as Ethereum, Arbitrum, Base, Optimism, and ZKsync Era, thus leading to the new phenomenon of inscriptions. They enable users to embed arbitrary data directly onto these blockchains. This process involves attaching data, such as text, images, or code, to transactions or smart contracts, effectively “inscribing” this information permanently into the blockchain. Users quickly started to use inscriptions to mint and trade tokens, instead of using the already available smart contract functionality. Similarly to what happened with Bitcoin ordinals, inscriptions triggered a significant increase in Transaction-per-Second (TPS)on these chains, reaching yearly transaction highs in December 2023. During this time, inscription transactions represented over $80$% of all network transactions [1]. Due to their apparent popularity, various marketplaces dedicated to the trading of ordinals and inscriptions began to pop-up [8, 17, 48]. In particular, even the leading crypto-exchange, Binance, supports ordinals and inscriptions [9].
During the inscription boom at the break of 2023 and early 2024, Zero Knowledge (ZK)rollups — a type of L2blockchain — witnessed notably lower gas costs compared to Ethereum, increasing the interest of users, as the cost of participating in the inscription boom became considerably more affordable compared to Bitcoin and Ethereum. However, some rollups experienced downtime of approximately $78$ minutes [16, 36]. On March 13, 2024, Ethereum underwent the Dencun upgrade [2], which introduced blobs as temporary data storage. Blobs decreased gas costs for Ethereum rollups, but also raised concerns about the impact on inscription-based tokens ownership, since they rely on out-of-the-blockchain applications and therefore presumably assume data availability [27]. Note that after the 18-day period, blobs may disappear from the Ethereum blockchain, as they are designed to support state verification for rollups. Consequently, inscription-related data might only be available off-chain in the indexer of their platform creator if the rollup chooses not to store all their transaction data.
We summarize the contributions of this paper as follows.
Causes of transaction surge at the end of 2023.  Our analysis reveals that the surge in transaction volume were primarily driven by the minting of new inscription-based meme coins. These tokens were later traded for a fraction of the costs compared to inscriptions on the Bitcoin or Ethereum networks, with prices in the range of then cents to one US dollar. We found that the inscription standards on these chains are based on the BRC-20 standards with some modifications to facilitate trading [35, 40, 46] and listing on NFTmarketplaces [8, 17, 48]. Specifically, on December 17, 2023, for several hours, over 96% of all transactions on ZKsync Era involved the minting of ZRC-20 sync tokens. Other blockchains also exhibited a high percentage of inscription-related transactions. Arbitrum recorded the highest percentage, with over 89%, followed by Ethereum at over 53%, Base at over 37%, and Optimism at the lowest, with over 35%.
Impact on EVMblockchain performance and cost efficiency.  Given that certain L2 solutions experienced downtime, and a significant portion of transactions during late 2023 and early 2024 were inscriptions, we seek to evaluate the performance of these L2 scaling solutions during this stress-test. We found that the increased transaction activity resulted in a decrease in gas fees, positively affecting the users of the rollup chains. With more L2 transactions compressed into a single transaction batch, the median costs of L2 transactions decreased. Unlike Ethereum, where gas fees increased during periods of transaction surges, median transaction fee remained the same on Base and Optimism, and decreased on Arbitrum and ZKsync Era. Additionally, ZKsync Era, as a ZK rollup, demonstrated a stronger reduction in fees than Arbitrum, an optimistic rollup.
Speculative purchases.  We observed that across all chains, most on-chain inscription actions were for minting inscriptions rather than selling them. For example, the percentage of inscription minting activities were as follows: Arbitrum at 99.49%, Base at 99.52%, Ethereum at 98.71%, Optimism at 99.73%, and ZKsync Era at 99.6%. On ZKsync Era, the minting process was completed for three tokens: zrc-20 sync, era-20 bgnt and layer2-20 \$L2, which are traded on ZKsync Era for a fraction of the costs compared to the Ethereum and Bitcoin networks. Even with observed prices in the range of \$0.1 to \$1, the costs of minting and listing are covered for the seller. The trading activity for the era-20 bgnt token continued after the initial inscription boom.
Scientific reproducibility.  Gathering and processing blockchain data can be costly both financially and computationally, even for a single chain. Analyzing multiple chains, as in this study, increases complexity because most of the data are not even easily accessible. To promote scientific reproducibility and enable other researchers and companies to study blockchain data and inscriptions, we plan to make our datasets and scripts publicly available.
2 Background on Rollups
Scalability challenges and gas price fluctuations in blockchain networks are typically addressed through Layer-2 (L2)solution. While, L1is the foundational layer for blockchains, L2is a secondary layer that provides scalability and relies on L1for security. This also introduces rollups, a type of non-custodial L2blockchain, and provides the necessary background for this paper.
Layer-1 (L1).  This approach involves creating an entirely new blockchain with unique consensus mechanisms and dedicated physical infrastructure to maintain network consensus. Notable L1blockchains include Bitcoin, Ethereum, Solana, Avalanche, and Polkadot. L1blockchains typically prioritize decentralization and security, often at the expense of scalability. To address this limitation, L2-rollup solutions were proposed to help scale L1blockchains.
Layer-2 (L2).  In contrast, L2scaling adopts a different strategy. It involves executing computations outside the main (L1) blockchain and then recording the results or final state on the underlying chain. Major types of L2solutions include state (payment) channels, plasma, and rollups. Plasma and state channels focus on moving both data and computation off-chain. Rollups, on the other hand, move computation and state storage off-chain while retaining compressed data for each transaction on the underlying chain.
Rollups.  They function as a form of L2scaling by performing computations off-chain and storing the results on the L1chain. They operate like blockchains, generating blocks and then recording batches of these blocks on the L1chain. However, in the rollup environment, operators are not inherently trusted. This lack of trust means that operators could potentially behave maliciously by halting block production, generating invalid blocks, withholding data, or engaging in other adversarial activities [10]. To mitigate these risks, rollups typically implement robust security measures to ensure the integrity and security of the system. In that regard, ensuring the correctness of the state in rollups involves two main approaches: optimistic and Zero Knowledge Proof (ZKP). Some rollups, such as ZKsync Era, employ ZKP, which involve computing cryptographic proofs to validate the correctness of the computations. On the other hand, rollups like Arbitrum and Optimism adopt an optimistic approach, which assumes all statements are true unless someone in the network contests them.
Sequencers.  These entities are responsible for aggregating, ordering, and rolling up the transactions to the L1chain. By bundling transactions, rollups effectively achieve cost savings in terms of gas fees.
Verifiers.  Operating as smart contracts on the L1chain, verifiers play a crucial role in validating the transactions stored by the sequencer. They ensure the correctness of the transactions, thereby maintaining the integrity of the overall system [15, 34, 45].
3 Inscriptions and Ordinals
@@
Inscriptions and ordinals involve the recording of arbitrary data on the blockchain. On EVM-compatible chains, users encode HEX data into the transaction input call data (refer to column JSON in Table 1), usually also setting the transaction’s from and to attributes with the same user addresses. This structure constitutes a self-transfer made from a user to their own address.
@@ 3.1 Operation Types
There are various protocol standards that define the structure inscriptions. BRC-20 was the first standard proposed for Bitcoin, and includes three main types of operations: deploy, mint, and transfer, each encoded in a single transaction [39]. The protocol standards on EVM-compatible chains extend this set of operations by adding, for example, claim, list, buy, sell, among others.
Table 1: Example of inscriptions data recorded on-chain: protocol name, operation, token name (Tick), total supply, and maximum amount of tokens minted each round (Limit)
@@ 3.2 Comparison with NFTs and ERC-20s
However, due to their lack of smart contract support, inscription-based tokens offer limited functionality. They are primarily utilized for speculative purposes and fall into the category of meme-coins [39]. Based on the inscriptions technology, it is possible to mint a single NFT-style token or a group of tokens with a predefined token supply in circulation. Nevertheless, it is the off-chain indexer, managed by the token’s creator, responsible for ensuring the data integrity of minted tokens. Trading of inscription-based tokens predominantly occurs within inscriptions marketplaces since they are not compatible with the ERC-20 standard required by DEXs [44].
@@ 4 Data Collection
We then collected blockchain data from various sources. For ZKsync, we use data obtained from an archive node and a Web3-compatible API for ZKsync Era. For Arbitrum, Base, Ethereum, and Optimism, we use data sources from Nansen BigQuery [26]. These data sources enable us to collect comprehensive data containing all information recorded on the evaluated chains. This includes data about blocks, transactions, and events (or logs triggered during the execution of smart contracts) specifically related to inscriptions. Table 2 provides an overview of the specifics of our dataset.
@@
ZKsync Era $481687$ $35.62$ $168.05$ $3$ $1$ $40770$
The distribution of inscription transactions per address is highly skewed towards a few participants, as shown in Table 3. Most inscription issuers initiated at least three transactions related to inscriptions. Arbitrum exhibited the highest median, with at least six transactions per issuer. Table 3 also highlights that a small number of issuers were responsible for a significant fraction of the total inscriptions. For instance, a single issuer accounted for $67713$ transactions to claim their inscriptions, followed by ZKsync Era with $40770$ transactions.
@@ 5.3 Inscription Trading
Table 4: Inscription operations cost in GWei for major token.
@@ 6 Discussion
Inscriptions: good or bad?  A report by Binance on BRC-20 highlights that the unexpected surge in transactions resulted in an increase in transaction fees for many blockchains [35]. The report suggests that such an increase in fees can be seen as a natural progression in blockchain adoption. Therefore, attribution of this spike solely to BRC-20s is considered irrelevant. More concerning, however, is that other protocols, such as Arbitrum, experienced downtime of approximately 78 minutes [36], while ZKsync Era saw an increase in TPS, with nearly 96% of all transactions being related to inscriptions over the course of several hours. Thus, causing users to worry about delayed transaction inclusion. This sparked a broader debate within the blockchain community about whether ordinals and inscriptions truly benefit users.
@@
Hope or Hype?  During minting periods, platforms utilized various channels to engage users, including communities, influencers, and specific chains that amplified them. Consequently, numerous users issued inscriptions across different platforms [8, 17, 48]. Understanding user expectations, often influenced by FOMO, regarding inscriptions is crucial. For instance, are inscriptions more about hope or hype? This question relates to whether inscriptions are a novel concept or just a passing trend. In the case of Bitcoin, there was a significant surge in interest and user participation during the BRC-20 movement, followed by a subsequent decline in enthusiasm over the following months [39, 40]. Our work shows similar results for EVM-compatible chains: inscriptions peak initially but witness a decline in user interest over time, as shown in Section 5.1. The evolution of inscriptions fuels discussions on innovative features and integration with tokens and NFTs. Despite potential profitability, marketplace platforms encounter challenges in translating inscriptions into substantial profits for users, particularly across different chains. Due to the nature of inscriptions, which rely on off-chain applications, there is currently no mechanism to prevent users from purchasing previously sold inscriptions. Additionally, users can utilize software programs to automatically claim the majority of inscriptions during the minting or claiming period. Hence, should inscriptions be distributed directly to wallets, similar to traditional airdrop methods [24]?
@@ 8 Conclusion
First, we observed that, on specific peak days, transactions related to inscriptions accounted for nearly 90% on Arbitrum and ZKsync Era, and over 53% on Ethereum, and almost 35% on Optimism and Base. The surge in on-chain inscriptions overwhelmed some rollups, causing them instability due to the increased transaction volume. This resulted in downtimes and delays in transaction processing.
@@
Finally, to promote scientific reproducibility and facilitate further research, we plan to make our datasets and scripts publicly available.
Acknowledgments
This research article is a work of scholarship and reflects the authors’ own views and opinions. It does not necessarily reflect the views or opinions of any other person or organization, including the authors’ employer. Readers should not rely on this article for making strategic or commercial decisions, and the authors are not responsible for any losses that may result from such use.

### v3

@@ Abstract
This paper examines inscription-related transactions on Ethereum and major EVM-compatible rollups, assessing their impact on scalability during transaction surges. Our results show that, on certain days, inscriptions accounted for nearly 90% of transactions on Arbitrum and ZKsync Era, while 53% on Ethereum, with 99% of these inscriptions involving meme coin minting. Furthermore, we show that ZKsync and Arbitrum saw lower median gas fees during these surges. ZKsync Era, a ZK-rollup, showed a greater fee reduction than the optimistic rollups studied—Arbitrum, Base, and Optimism.
@@ 1 Introduction
Layer-2 (L2)scaling solutions have emerged as a response to the scalability challenges of Layer-1 (L1)blockchains such as Bitcoin and Ethereum. These solutions aim to enhance transaction throughput and cost efficiency while inheriting the security guarantees of their underlying L1blockchain. L2sexecute transactions off-chain and subsequently settling state transitions on an L1blockchain.
Despite their increasing adoption, most L2solutions had not undergone significant stress testing until late 2023, when an unprecedented spike in transaction volume occurred. This surge was largely driven by the emergence of inscriptions, a phenomenon inspired by Bitcoin ordinals [27]. Ordinals, introduced in March 2023 [5], enable the assignment of unique identifiers to individual satoshis, effectively transforming them into Non-Fungible Tokens (NFTs). This concept extended beyond Bitcoin, reaching Ethereum and its major L2rollups, where inscriptions allowed users to embed arbitrary data, such as text, images, or code, directly onto the blockchain. By late 2023, inscription transactions accounted for over 80% of all transactions on some Ethereum Virtual Machine (EVM)chains [1], significantly influencing network performance, transaction costs, and user behavior.
This rapid increase in transaction volume led to temporary service disruptions, with certain rollups experiencing downtime of approximately 78 minutes [12, 23]. However, during this surge, Zero Knowledge (ZK)rollups—one type of L2blockchain—offered notably lower gas costs compared to Ethereum, making them an attractive alternative for inscription trading. Furthermore, Ethereum’s Dencun upgrade in March 2024 introduced blobs as temporary data storage, reducing gas fees for rollups but raising concerns about the long-term availability of inscription-related data [20]. Since blobs are designed for ephemeral state verification, inscription-related data may not persist on-chain, relying instead on off-chain indexers managed by platform creators.
In this paper, we investigate the performance and user behavior of EVM-compatible chains during the inscription boom. Our study focuses on Ethereum and its four major rollups—Arbitrum, Base, Optimism, and ZKsync Era. We address the following research questions:
What caused the transaction surge on EVMchains at the end of 2023?  We analyze transaction data of Ethereum and its four major rollups. We identified that these transactions were due to inscription-based meme coins and were minted and traded at significantly lower costs compared to Bitcoin ordinals or Ethereum inscriptions. Our findings indicate that these inscriptions adopted modified versions of the BRC-20 standard [22, 27, 30], facilitating their integration into NFT marketplaces [7, 13, 32]. Specifically, on December 17, 2023, over 88% of all transactions on ZKsync Era were due to inscriptions. Other blockchains also exhibited a high percentage of inscription-related transactions. Arbitrum recorded the highest percentage, with over 89%, followed by Ethereum at over 53%, Base at over 37%, and Optimism, with over 35% of the total inscription transactions.
What impact did inscription transactions have on the cost and performance of EVMblockchains?  We evaluate how inscription-driven transaction spikes affected network stability, gas fees, and transaction throughput. Our analysis shows that, unlike Ethereum, where transaction surges led to increased gas fees, some rollups experienced reduced median transaction costs. We also compare the cost efficiency of ZK rollups versus optimistic rollups.
Why did users choose inscription-based meme coins over ERC-20 tokens?  We examine the user interaction with inscriptions, finding that inscription transactions predominantly involved minting rather than trading. Across all analyzed chains, over 99% of inscription-related transactions were for minting activities, with some tokens, such as zrc-20 sync, sustaining trading activity beyond the initial boom. For example, in comparison to the overall number of inscription transactions, the percentage of inscription minting or claiming activities was as follows: Arbitrum at 99.82%, Base at 99.99%, Ethereum at 99.92%, Optimism at 99.98%, and ZKsync Era at 99.91%.
To promote scientific reproducibility, we make our datasets and analysis scripts publicly available [17]. By providing insights into the inscription-driven stress test of EVMchains, our study contributes to the broader understanding of blockchain scalability, economic incentives, and data availability in the evolving blockchain ecosystem.
2 Background on Inscriptions
@@
Inscriptions involve the recording of arbitrary data on the blockchain. On EVM-compatible chains, users encode HEX data into the transaction input call data (refer to column JSON in Table 1), usually also setting the transaction’s from and to attributes with the same user addresses. This structure constitutes a self-transfer made from a user to their own address.
@@ 2.1 Operation Types
There are various protocol standards that define the structure of inscriptions. BRC-20 was the first standard proposed for Bitcoin, and includes three main types of operations: deploy, mint, and transfer, each encoded in a single transaction [26]. The protocol standards on EVM-compatible chains extend this set of operations by adding, for example, claim, list, buy, sell, among others.
Table 1: Example of inscriptions data recorded on-chain: protocol name, operation, token name (Tick), total supply, and maximum amount of tokens minted each round (Limit).
@@ 2.2 Comparison with NFTs and ERC-20s
However, due to their lack of smart contract support, inscription-based tokens offer limited functionality. They are primarily utilized for speculative purposes and fall into the category of meme-coins [26]. Based on the inscriptions technology, it is possible to mint a single NFT-style token or a group of tokens with a predefined token supply in circulation. Nevertheless, it is the off-chain indexer, managed by the token’s creator, responsible for ensuring the data integrity of minted tokens. Trading of inscription-based tokens predominantly occurs within inscriptions marketplaces since they are not compatible with the ERC-20 standard required by Decentralized Exchanges (DEXs) [29].
@@ 3 Data Collection
We then collected blockchain data from various sources and made them and our scripts available in a public repository [17]. For ZKsync, we use data obtained from their official archive node and a Web3-compatible API for ZKsync Era [9]. For Arbitrum, Base, Ethereum, and Optimism, we use data sources from Nansen BigQuery [19]. These data sources enable us to collect comprehensive data containing all information recorded on the evaluated chains. This includes data about blocks, transactions, and events (or logs triggered during the execution of smart contracts) specifically related to inscriptions. Table 2 provides an overview of the specifics of our dataset.
@@
ZKsync Era $481687$ $35.62$ $245.35$ $3$ $1$ $40770$
The distribution of inscription transactions per address is highly skewed towards a few participants, as shown in Table 3. Most inscription issuers initiated at least three transactions related to inscriptions. Arbitrum exhibited the highest median, with at least six transactions per issuer. Table 3 also highlights that a small number of issuers were responsible for a significant fraction of the total inscriptions. For instance, a single issuer accounted for $67713$ transactions to claim their inscriptions, followed by ZKsync Era with $40770$ transactions. These are likely bots accounts.
@@ 4.3 Inscription Trading
Table 4: Inscription operations cost in GWei for major tokens.
@@ 5 Discussion
Inscriptions: good or bad?  A report by Binance on BRC-20 highlights that the unexpected surge in transactions resulted in an increase in transaction fees for many blockchains [22]. The report suggests that such an increase in fees can be seen as a natural progression in blockchain adoption. Therefore, attribution of this spike solely to BRC-20s is considered irrelevant. More concerning, however, is that other protocols, such as Arbitrum, experienced downtime of approximately 78 minutes [23], while ZKsync Era saw an increase in Transaction-per-Second (TPS), with nearly 96% of all transactions being related to inscriptions over the course of several hours. Thus, causing users to worry about delayed transaction inclusion. This sparked a broader debate within the blockchain community about whether ordinals and inscriptions truly benefit users.
@@
Hope or Hype?  During minting periods, platforms utilized various channels to engage users, including communities, influencers, and specific chains that amplified them. Consequently, numerous users issued inscriptions across different platforms [7, 13, 32]. Understanding user expectations, often influenced by Fear-Of-Missing-Out (FOMO), regarding inscriptions is crucial. For instance, are inscriptions more about hope or hype? This question relates to whether inscriptions are a novel concept or just a passing trend. In the case of Bitcoin, there was a significant surge in interest and user participation during the BRC-20 movement, followed by a subsequent decline in enthusiasm over the following months [26, 27]. Our work shows similar results for EVM-compatible chains: inscriptions peak initially but witness a decline in user interest over time, as shown in Section 4.1. The evolution of inscriptions fuels discussions on innovative features and integration with tokens and NFTs. Despite potential profitability, marketplace platforms encounter challenges in translating inscriptions into substantial profits for users, particularly across different chains. Due to the nature of inscriptions, which rely on off-chain applications, there is currently no mechanism to prevent users from purchasing previously sold inscriptions. Additionally, users can utilize software programs to automatically claim the majority of inscriptions during the minting or claiming period. Hence, should inscriptions be distributed directly to wallets, similar to traditional airdrop methods [18]?
@@ 7 Conclusion
First, we observed that, on specific peak days, transactions related to inscriptions accounted for nearly 90% on Arbitrum and ZKsync Era, and over 53% on Ethereum, and almost 35% on Optimism and Base. The surge in on-chain inscriptions overwhelmed some rollups, causing their instability due to the increased transaction volume. This resulted in downtimes and delays in transaction processing.
@@
Finally, to promote scientific reproducibility and facilitate further research, we made our data set and scripts publicly available [17].